OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!!! HURRY PLEASE! What is the slope of a line parallel to the line 2x - 3y = 6? A. 2/3 B. 4/3 C. 1/5 D. 1

OpenStudy (anonymous):

A I hope I helped

OpenStudy (pawanyadav):

Parallel lines have same slope

OpenStudy (anonymous):

yea i go with A.

OpenStudy (anonymous):

set it to y intercept format to be able to more easily find slope. y=mx+b; m=slope, 2x-3y=6 -3y=-2x+6 y=2/3x-2 A line has the SAME slope as the other line(s) that it is parallel to. and in this case we know that slope (m) is equal to 2/3